# Which Augmentation Should I Use? An Empirical Investigation of Augmentations for Self-Supervised Phonocardiogram Representation Learning

### v3

@@ Abstract
Despite the recent increase in research activity, deep-learning models have not yet been widely accepted in several real-world settings, such as medicine. The shortage of high-quality annotated data often hinders the development of robust and generalizable models, which do not suffer from degraded effectiveness when presented with newly-collected, out-of-distribution (OOD) datasets. Contrastive Self-Supervised Learning (SSL) offers a potential solution to labeled data scarcity, as it takes advantage of unlabeled data to increase model effectiveness and robustness. However, the selection of appropriate transformations during the learning process is not a trivial task and often leads to degraded performance or even breaks down the ability of the network to extract meaningful information. In this research, we propose uncovering the optimal augmentations for applying contrastive learning, detecting abnormalities in 1D phonocardiogram (PCG) samples and thus learning a generalized representation of the signal. Specifically, we perform an extensive comparative evaluation of a wide range of audio-based augmentations, evaluate trained classifiers on multiple datasets across different downstream tasks, and finally report on the impact (positive and negative) of each augmentation in model training. We experimentally demonstrate that, depending on its training distribution, the effectiveness of a fully-supervised model can degrade up to 32% when evaluated on unseen data, while SSL models only lose up to 10% or even improve in some cases. We argue and experimentally demonstrate that, contrastive SSL pretraining can assist in providing robust classifiers which can generalize to unseen, OOD data, without relying on time- and labor-intensive annotation processes by medical experts. Furthermore, the proposed extensive evaluation protocol sheds light on the most promising and appropriate augmentations for robust PCG signal processing, by calculating their effect size on model training. Finally, we provide researchers and practitioners with a roadmap towards producing robust models for PCG classification, in addition to an open-source codebase for developing novel approaches.
keywords:
Self-Supervised learning, Contrastive learning, Deep learning, Phonocardiogram classification, OOD representation learning
@@
[inst1]organization=Department of Informatics and Telematics, Harokopio University of Athens,addressline=Omirou 9, city=Tavros, postcode=177 78, state=Athens, country=Greece
[…]
[inst2]organization=Department of Biosciences and Nutrition, Karolinska Institute,addressline=Blickagangen 16, city=Flemingsberg, postcode=141 52, state=Stockholm, country=Sweden
1 Introduction
@@
Abnormal phonocardiogram (PCG) signals are often an important indicator of heart diseases or deficiencies linked to increased mortality gbd_2017_causes_of_death_collaborators_global_2018 . The detection of such indicators in a non-invasive, accurate, and robust manner could assist towards prevention or timely treatment of such diseases, especially when access to healthcare services is limited.
@@
Automated methods for detecting abnormal samples commonly rely on audio recordings; typically, audio is processed and then used to train classifiers that detect abnormal (i.e., non-healthy) samples and sometimes discriminate between different diseases (e.g., khan2020 ). However, the audio quality of collected signals can be affected by several circumstances. While typical audio-related factors include microphone type or sampling rate, among others, for the case of PCG recordings additional factors can significantly affect the quality of their recording, such as placement of the microphone jaros2023novel ; fontecave2019 , body size, skin, noise giordano2021 ; huisa2023 , etc. The above cases can prove even more challenging, such as when screening an unborn child/fetus, where PCG can only be captured through the mother’s belly kahankova2023 . In addition to the quality of the recording conditions, the training of accurate ML models is heavily dependent on annotated data, the collection of which proves highly time- and cost-consuming. The above is especially important in the biomedical domain, where available samples are scarce to begin with.
As a result, classification models that have been trained on data originating from a single source or on low quality signals fail to generalize to previously unseen data distributions, which do not adhere to the i.i.d. assumption (i.e they are not independently and identically distributed), leading to degraded effectiveness recht2019imagenet ; 10233054 . In this work, we propose training robust classifiers for PCG signal classification by employing contrastive self-supervised learning (SSL) and investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets. However, as the effectiveness of an backbone encoder trained via SSL heavily depends on the adopted augmentation policy Reed_2021_CVPR , we also research the appropriate transformations to be applied on PCG signals and report their effect in model training. Specifically, we attempt to provide insight into the following research questions:
  • 1.
@@
    Does contrastive SSL prove a valid training method for extracting robust and meaningful PCG representations?
  • 2.
@@
    If so, which augmentations or transformations lead to such representations, proving the most effective, and which actually inhibit training?
  • 3.
@@
    Is a classifier trained on the above representations able to maintain adequate effectiveness and demonstrate generalization capabilities when evaluated on OOD data, in contrast to its fully supervised counterpart?
@@
To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets, (b) examine the effectiveness on a global “normal vs. abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable), (c) evaluate all trained classifiers on OOD data that are left-out during training, and finally, (d) calculate the effect size of each augmentation. To the best of our knowledge, this is the first paper researching the effect of each applied augmentation on learning robust OOD PCG representations. To facilitate further research in the field and provide a comprehensive baseline, we have open-sourced our codebase [^1] Code available at: https://github.com/aristotelisballas/listen2yourheart.
@@
The rest of the paper is organized as follows. Initially, section 2 presents previous work on PCG-based classification, SSL training, and model robustness. Followingly, section 3 presents the methodology, including training approach, network architecture, augmentations, and datasets. Finally, section 4 presents the evaluation framework and results, while section 5 concludes the paper.
2 Related Work
@@
Accurate and robust PCG classification has major implications in the well-being of patients since it can often be an indicator of heart deficiencies. However, the limited availability of annotated data has hindered the progress of research in the field, along with biosignal classification in general. In this section, we focus on the most important papers and previous works in DL for PCG classification and in SSL for 1D biosignal processing. We also refer to methods which have been specifically developed for robust 1D biosignal classification.
2.1 Deep Learning for Phonocardiogram Classification
@@
The release of publicly available PCG datasets under the scope of the 2016 liu_open_2016 and 2022 reyna_heart_2023 PhysioNet goldberger_physiobank_2000 challenges, has a sparked a line of research in deep learning for PCG classification. The majority of previous works have mainly proposed either 2D or 1D CNN architectures for tackling this particular problem. The main difference lies in whether the proposed models are trained directly on the one dimensional recorded auscultation waves or on spectral images extracted from the audio recordings. For example, multiple recent papers lu2022lightweight ; 10081740 ; 10081866 ; 10081768 propose combining mel-spectrogram images extracted from PCG signals with encoded patient demographic data, in a dual-branch CNN model. The authors find that the combination of the above representations leads to improved accuracy in the heart murmur detection task. Similarly, in another work utilizing frequency domain features, the authors of 10081900 design a hierarchical deep CNN architecture that extracts and combines features from multiple-scales of PCG Mel-Spectrograms. A more recent work, 10122565 , proposes training an ensemble of 15 ResNets He_2016_CVPR with channel-wise attention mechanisms on similar spectrograms, for grading murmur intensity in PCG recordings.
@@
Although few, 1D models also prove promising in the heart-murmur detection domain. In alkhodari_ensemble_2022 the authors employ a transformer-based neural network and train it on 1D wavelet power features extracted from raw PCG signals. On the other hand, 10081735 proposes using a U-Net to predict heart states from the raw PCG signal and combine the predictions with a ResNet model for the final classification. In a different approach, the authors of 10081890 suggest that the combination of a 1D CNN-Bi-LSTM model coupled with attention mechanisms, trained in conjunction with a feature-based classifier leads to robust abnormal PCG signal detection.
@@
In this work, we build upon our preliminary research ballas_listen2yourheart_2022 for the 2022 PhysioNet challenge. In particular, we demonstrate the value of adopting self-supervised contrastive learning for robust PCG classification and also report the top augmentations which should be implemented in such a framework. To our knowledge, this is the first study extensively researching the benefits of each augmentation, and their combinations, for contrastive SSL in 1D PCG signal processing. The aim of this study is to provide future researchers with best-practices regarding PCG signal augmentation, while also propose a structured protocol for the evaluation of developed methods.
@@
Figure 1: Illustration of the proposed experiment pipeline for training and evaluating the effectiveness and robustness of a model trained via Self-Supervised Contrastive Learning for PCG classification. The framework has six steps. In the first step, all datasets are prepared and homogenized into a common format, as described in Section 3.1. In the second step, each signal is augmented to two versions. In the next step, the unlabeled and augmented signals are used to train the backbone encoder (Fig. 2 left) to maximize the agreement between representations originating from the same initial signal. Following the pretraining phase, in step 4, a classification head (Fig. 2 right) is attached to the frozen weights of the pretrained encoder. The classifier weights are fine-tuned on data drawn from one of the Pascal, PhysioNet2016 or PhysioNet2022 datasets, in a fully supervised manner. The final model is then evaluated on the test split of the same dataset. In the fifth step of the framework, the generalization ability of the fine-tuned model is assesed on signals drawn from datasets which are left-out (”unseen”) during the training process in step 4. We argue that given sufficient data and appropriate augmentations, the pretrained encoder will be able to extract a generalized PCG representation regardless the dataset. Finally, we repeat the fine-tuning and OOD evaluation steps for each remaining dataset (step 6), to assess the robustness of our method when the classifier has been trained on different signal distributions.
2.2 Self-Supervised Learning in 1D biosignal classification
@@
Self-supervised representation learning (SSL) 9770283 was initially proposed for diminishing the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks misra2020self to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding oord_representation_2019 , the goal of Self-Supervised Contrastive Learning 9462394 ; kumar_contrastive_2022 is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions NEURIPS2021_fa14d4fe . The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which can take advantage of the small amount of data but also yield adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
@@
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection 9837871 . The authors of mohsenvand_contrastive_2020 propose augmenting multi-channel EEG signals by recombining several channels into additional channels which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying pre-defined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of banville_self-supervised_2019 argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in banville2021uncovering where contrastive predictive coding is implemented on similar SSL-learned features to both predict sleep-stages and detect pathologies in multi-channel EEG signals. Finally, kostas_bendr_2021 proposed BENDR, which combines attention mechanisms vaswani_attention_2017 and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
@@
Several SSL methods have been proposed for ECG signal analysis as well mehari_self-supervised_2022 , most of which focus on either emotion recognition or pathology detection. For example, in sarkar_self-supervised_2022 the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the art results. In another paper studying emotion recognition from ECG signals, vazquez2022transformer introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal kiyasseh_clocs_2021 , implementing attention mechanisms oh_lead-agnostic_2022 or combining wavelet transformations and random crops of the signal chen_clecg_2021 for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, lan_intra-inter_2022 propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
@@
Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models 9346072 . For example, liu_practical_2023 introduces NeuroPose, a system for 3D hand pose tracking which aims to alleviate the problems imposed due to different sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user. Similarly, the authors lai_contrastive_2022 employ a contrastive loss to pretrain a model on EMG signals for hang-gesture recognition.
2.3 Model Robustness in 1D biosignal classification

### v4

# Date of publication xxxx 00, $0$, date of current version xxxx 00, $0$. PREPRINT of Accepted Manusript

@@ Abstract
Despite recent advancements in deep learning, its application in real-world medical settings, such as phonocardiogram (PCG) classification, remains limited. A significant barrier is the lack of high-quality annotated datasets, which hampers the development of robust, generalizable models that can perform well on newly collected, out-of-distribution (OOD) data. Self-Supervised Learning (SSL), particularly contrastive learning, has shown promise in mitigating the issue of data scarcity by leveraging unlabeled data to enhance model robustness and effectiveness. Even though SSL methods have been proposed and researched in other domains, works focusing on the impact of data augmentations on model robustness for PCG classification is limited. In particular, while augmentations are a key component in SSL, selecting the most suitable transformations during the training process is highly challenging and time-consuming. Improper augmentations can lead to substantial performance degradation, even hindering the network’s ability to learn meaningful representations. Addressing this gap, our research aims to explore and evaluate a wide range of audio-based augmentations and uncover combinations that enhance SSL model performance in PCG classification. We conduct a comprehensive comparative analysis across multiple datasets and downstream tasks, assessing the impact of various augmentations on model performance and generalization. Our findings reveal that depending on the training distribution, augmentation choice significantly influences model robustness, with fully-supervised models experiencing up to a 32% drop in effectiveness when applied to unseen data, while SSL models demonstrate greater resilience, losing only 10% or even improving in some cases. This study also sheds light on the most promising and appropriate augmentations for robust PCG signal processing, by calculating their effect size on model training. These insights equip researchers and practitioners with valuable guidelines for building more robust, reliable models in PCG signal processing.
Index Terms:
Contrastive learning, deep learning, OOD representation learning, phonocardiogram classification, self-supervised learning
@@
=-21pt
I Introduction
@@
Abnormal phonocardiogram (PCG) signals are often regarded as an important indicator of heart diseases or deficiencies linked to increased mortality [1]. Since the collection of PCG signals is non-invasive, the automatic, accurate detection of abnormalities in a sample has the potential to assist in early prevention or timely treatment of such diseases, especially when access to healthcare services for certain populations is limited [2].
@@
The efficacy of automated methods for detecting abnormal audio samples, depends to a large extent on the quality and nature of the audio recordings themselves. Typically, signals are first manually analyzed and processed before being used to train classifiers that discriminate between normal recordings and samples that include indicators for different diseases [3]. However, these classifiers often fail to produce generalizable and robust representations of signals due to distribution shifts present in either the training data or on the data that the model is evaluated on after training [4]. For example, a robust classifier should be able to maintain its accuracy on signals that have been collected in different environments, from multiple devices, and with varying sampling rates. Specifically for the case of PCG recordings, additional factors such as microphone placement [5, 6], body size, and skin type [7, 8], among others, can significantly affect the signals. The above cases prove even more challenging when screening an unborn child/fetus, where PCG can only be captured through the mother’s abdomen [9]. In addition to recording conditions, the training of accurate ML models is dependent on annotated data, the collection of which proves highly time- and cost-consuming. The above is especially important in the biomedical domain, where available samples are scarce to begin with. As a result, classification models that have been trained on data originating from a single source or on low-quality signals, fail to generalize to previously unseen data distributions, which do not adhere to the i.i.d. assumption (i.e., are not independently and identically distributed), leading to degraded effectiveness [10, 4].
A possible solution to the above limitations, which has proven advantageous in various ML fields, is the utilization of unlabeled data [11, 12]. By training models on a large corpus of unlabeled data via self-supervised learning (SSL), the feature extractor is able to learn general representations and avoid overfitting on its training distributions. Nonetheless, even though there have been works researching the benefits of SSL in biosignal analysis, they are quite limited and primarily focus on either electrocardiography, electromyography or electroencephalography signals (see Section II-B). Therefore, in this work we turn our attention to the Phonocardiogram domain and propose training robust classifiers for PCG signal classification via contrastive self-supervised learning, leveraging both labeled and unlabeled datasets. Furthermore, as the effectiveness of a backbone encoder trained via SSL heavily depends on the adopted augmentation policy [13]. The design of a fitting augmentation policy for each signal distribution is arguably the most critical and time-consuming step in an SSL framework. To facilitate future research in this domain, we perform an exhaustive evaluation on a large number of transformation combinations (150150~{}150150) and models (40004000~{}40004000) applied on PCG signals and report their effect on training and generalization. Specifically, we attempt to provide insight into the following research questions:
@@
  • Does the application of contrastive SSL prove a valid training method for extracting robust and meaningful representations in the PCG classification domain?
@@
  • If so, which augmentations or transformations lead to such representations and prove the most effective, and which actually inhibit training?
@@
  • Is a classifier trained on the above representations able to maintain adequate effectiveness and demonstrate domain generalization capabilities when evaluated on OOD data, in contrast to its fully supervised counterpart?
@@
To answer the above research questions, we develop and implement an extensive comparative evaluation in which we: (a) investigate multiple combinations of augmentations, leveraging both labeled and unlabeled datasets; (b) examine the effectiveness on a global “normal vs. abnormal” downstream task, as well as downstream tasks specific to each dataset (where applicable); (c) evaluate all trained classifiers on OOD data that are left out during training; and finally, (d) calculate the effect size of each augmentation. To the best of our knowledge, this is the first paper to research the effect of each applied augmentation on learning robust OOD PCG representations. To facilitate further research in the field and provide a comprehensive baseline, we have open-sourced our codebase [^1] Code available at: https://github.com/aristotelisballas/listen2yourheart.
@@
The rest of the paper is structured as follows. Initially, Section II presents previous work on PCG-based classification, SSL training, and model robustness. Subsequently, Section III presents the methodology, including training approach, network architecture, augmentations, and datasets. Finally, Section IV presents the evaluation framework and results, while Section V concludes the paper.
II Related Work
@@
Accurate and robust PCG classification has major implications in the well-being of patients as it can often be an indicator of heart deficiencies. However, the limited availability of annotated and quality data has hindered the progress of research in the field, along with biosignal classification in general. In this section, we focus on the most important papers and previous works in DL for PCG classification and in SSL for 1D biosignal processing. We also refer to methods which have been specifically developed for robust 1D biosignal classification.
II-A Deep Learning for Phonocardiogram Classification
@@
The release of publicly available PCG datasets under the scope of the 2016 [14] and 2022 [15] PhysioNet [16] challenges, has a sparked a line of research in deep learning for PCG classification. The majority of previous works have mainly proposed either 2D or 1D CNN architectures for tackling this particular problem. The main difference lies in whether the proposed models are trained directly on the one dimensional recorded auscultation waves or on spectral images extracted from the audio recordings. For example, multiple recent papers [17, 18, 19, 20] propose combining mel-spectrogram images extracted from PCG signals with encoded patient demographic data, in a dual-branch CNN model. The authors find that the combination of the above representations leads to improved accuracy in the heart murmur detection task. Similarly, in another work utilizing frequency domain features, the authors of [21] design a hierarchical deep CNN architecture that extracts and combines features from multiple-scales of PCG Mel-Spectrograms. A more recent work, [22], proposes training an ensemble of 15 ResNets [23] with channel-wise attention mechanisms on similar spectrograms, for grading murmur intensity in PCG recordings.
@@
Although few, 1D models also prove promising in the heart-murmur detection domain. In [24] the authors employ a transformer-based neural network and train it on 1D wavelet power features extracted from raw PCG signals. On the other hand, [25] proposes using a U-Net to predict heart states from the raw PCG signal and combine the predictions with a ResNet model for the final classification. In a different approach, the authors of [26] suggest that the combination of a 1D CNN-Bi-LSTM model coupled with attention mechanisms, trained in conjunction with a feature-based classifier leads to robust abnormal PCG signal detection.
@@
In this work, we build upon our preliminary research [27] for the 2022 PhysioNet Challenge. In particular, we demonstrate the value of adopting self-supervised contrastive learning for robust PCG classification and also report the top augmentations that should be implemented in such a framework. Through our research, we aim to extensively investigate the benefits of each augmentation, and their combinations, for contrastive SSL in 1D PCG signal processing. The aim of this study is to provide future researchers with best practices regarding PCG signal augmentation, while also proposing a structured protocol for the evaluation of developed methods.
@@
Figure 1: Illustration of the proposed experiment pipeline for training and evaluating the effectiveness and robustness of a model trained via Self-Supervised Contrastive Learning for PCG classification. The framework has six steps. In the first step, all datasets are prepared and homogenized into a common format, as described in Section III-A. In the second step, each signal is augmented to two versions. In the next step, the unlabeled and augmented signals are used to train the backbone encoder (Fig. 2 left) to maximize the agreement between representations originating from the same initial signal. Following the pretraining phase, in step 4, a classification head (Fig. 2 right) is attached to the frozen weights of the pretrained encoder. The classifier weights are fine-tuned on data drawn from one of the Pascal, PhysioNet2016 or PhysioNet2022 datasets, in a fully supervised manner. The final model is then evaluated on the test split of the same dataset. In the fifth step of the framework, the generalization ability of the fine-tuned model is assessed on signals drawn from datasets which are left-out (”unseen”) during the training process in step 4. We argue that given sufficient data and appropriate augmentations, the pretrained encoder will be able to extract a generalized PCG representation regardless the dataset. Finally, we repeat the fine-tuning and OOD evaluation steps for each remaining dataset (step 6), to assess the robustness of our method when the classifier has been trained on different signal distributions.
II-B Self-Supervised Learning in 1D biosignal classification
@@
Self-supervised representation learning (SSL) [28] was initially proposed for mitigating the limit posed by the requirement of large annotated datasets for practical and production-ready deep learning systems. By taking advantage of common features present in unlabeled data, SSL approaches rely on pretext tasks [29] to exploit them and ultimately provide models which are able to extract generalized and robust representations. Inspired by advancements in contrastive predictive coding [30], the goal of Self-Supervised Contrastive Learning [31, 32] is to train a model to recognize augmented or transformed views of the same initial signal. By distinguishing between representations which originate from different original signals, the resulting model is able to extract invariant representations from its input data and reason upon features which remain stable across distinct data distributions [33]. The SSL paradigm seems to fit quite well in the medical and healthcare domain. In a field where sufficient annotated datasets are scarcely available, there proves a need to provide systems which capable of leveraging small amounts of data while yielding adequate and trustworthy inference results. To this end, several works have proposed SSL approaches for processing biosignals in the past, the majority of which have studied its advantages when extracting representations from either electroencephalography (EEG), electrocardiography (ECG) or electromyography (EMG) signals.
@@
Regarding EEG signals, several prior studies revolve around developing methods for emotion recognition and/or sleep stage detection [34]. The authors of [35] propose augmenting multi-channel EEG signals by recombining several channels into additional ones which represent different aspects of the EEG recording. On top of the combination, they augment each channel by applying predefined transformations and employ a contrastive loss for improved emotion recognition and sleep-stage classification. In another work, the authors of [36] argue that a model can learn informative EEG representations by predicting whether two windows of a signal are sampled from the same temporal context and use the above reasoning to design a relevant pretext task. The above study is extended in [37] where contrastive predictive coding is implemented on similar SSL-learned features to predict sleep stages and detect pathologies in multi-channel EEG signals. Finally, [38] proposed BENDR, which combines attention mechanisms [39] and contrastive learning to produce a model with increased generalization capabilities and fine-tune it for several downstream tasks in the EEG biosignal domain.
@@
Several SSL methods have been proposed for ECG signal analysis as well [40], most of which focus on either emotion recognition or pathology detection. For example, in [41] the authors argue that the training of a model to recognize different ECG signal transformations can lead to generalized and robust feature learning. They evaluate their method on four emotion recognition datasets and report state-of-the-art results. In another paper studying emotion recognition from ECG signals, [42] introduces a transformer model which is pretrained to predict masked values of the signal and then fine-tuned to predict the level of emotion in individuals. Regarding the pathology detection task, most prior works focus on discriminating normal ECG signals from others indicating a type of arrhythmia. Among the self-supervised approaches proposed for ECG pathology detection, contrastive learning is the most prevalent choice. Specifically, authors either propose exploiting the temporal and spatial invariances of the ECG signal [43], implementing attention mechanisms [44] or combining wavelet transformations and random crops of the signal [45] for the pretext contrastive task, before fine-tuning their resulting models for arrhythmia detection. In a different approach, [46] propose learning good representations by contrasting signals both at the inter- and intra-subject levels. Specifically, they argue that similarity should be maximized between signals which share a common label but are measured from different individuals.
@@
Finally, EMG signal processing is another interesting field which would greatly benefit from the application of robust prediction models [47]. For example, [48] introduces NeuroPose, a system for 3D hand pose tracking which aims to address challenges posed by varying sensor mounting or wrist positions, by pretraining an encoder-decoder model to recognize augmented views of a single user’s signal and then finetuning a downstream model to a subset of signals from a separate user. Similarly, the authors [49] employ a contrastive loss to pretrain a model on EMG signals for hand-gesture recognition.
II-C Model Robustness in 1D biosignal classification